# An analogue of Ramanujan’s identity for Bernoulli-Carlitz numbers

### v3

@@ 1. Introduction
(2.1) $e(z)$ $=e_{\tilde{\pi}R}(z)$
$=z\prod_{0\neq\lambda\in\tilde{\pi}R}\left(1-\frac{z}{\lambda}\right)$
$=z\prod_{0\neq A\in R}\left(1-\frac{z}{\tilde{\pi}A}\right)$
$=z\prod_{A\in R_{1}}\left(1-\left(\frac{z}{\tilde{\pi}A}\right)^{q-1}\right),$
@@
(2.2) $\frac{1}{e(z)}$ $=\frac{e^{\prime}(z)}{e(z)}$
$=\frac{1}{z}+\sum_{A\in R_{1}}\frac{-\frac{(q-1)z^{q-2}}{(\tilde{\pi}A)^{q-1}}}{1-\left(\frac{z}{\tilde{\pi}A}\right)^{q-1}}$
$=\frac{1}{z}-z^{q-2}\sum_{A\in R_{1}}\frac{1}{z^{q-1}-(\tilde{\pi}A)^{q-1}}.$
Thus for any $\alpha\in\Omega$ and $B\in R$, we have
(2.3) $\frac{1}{e(\alpha B)}=\frac{1}{\alpha B}-(\alpha B)^{q-2}\sum_{A\in R_{1}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}.$
@@
(2.4) $\quad\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j}\left((\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}\right)}$
$=\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j}}\frac{1}{(\alpha B)^{q-2}}\left(\frac{1}{\alpha B}-\frac{1}{e(\alpha B)}\right)$
@@
$\left.\qquad\qquad\qquad\qquad\qquad\quad-\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j+(q-2)}e(\alpha B)}\right).$
@@
(2.5) $\quad\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j}\left((\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}\right)}$
@@
$\quad-\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j+(q-2)}e(\alpha B)}$
@@
×$\quad\times\sum_{B\in R_{1}}\frac{B^{-(q-1)j-(q-2)}}{e(\alpha B)}.$
@@
(2.6) $\quad\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j}\left((\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}\right)}$
$=-\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{q-1}}$
$\quad+\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\alpha B)^{(q-1)(j-1)}(\tilde{\pi}A)^{q-1}\left((\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}\right)}$
$=-\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{q-1}}$
$\quad-\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)(j-1)}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)2}}$
$\quad+\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\alpha B)^{(q-1)(j-2)}(\tilde{\pi}A)^{(q-1)2}\left((\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}\right)}$
=$\cdots$= ⋯
[…]
$=-\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{q-1}}$
$\quad-\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)(j-1)}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)2}}$
[…]
$\quad-\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{q-1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}$
$\quad+\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}$
$=-\sum_{k=1}^{j}\left(\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)(j-k+1)}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)k}}\right)$
$\quad+\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}$
@@
$\quad+\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}},$
@@
(2.7) $\quad\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\alpha B)^{(q-1)(j+1)}\left((\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}\right)}$
@@
$\quad+\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}$
@@
$\quad+\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}.$
@@
(2.8) $\quad-\alpha^{-(\frac{q-1}{2})(j-1)-(q-2)}\sum_{B\in R_{1}}\frac{B^{-(q-1)j-(q-2)}}{e(\alpha B)}$
@@
$\quad+\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}.$
@@
(2.9) $\quad\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}$
$=\frac{1}{\tilde{\pi}^{(q-1)j}}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{A^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-\alpha^{\frac{q-1}{2}}\beta^{\frac{q-1}{2}}A^{q-1}}$
$=\frac{1}{\tilde{\pi}^{(q-1)j}}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{\alpha^{\frac{q-1}{2}}A^{(q-1)j}}\frac{1}{\alpha^{\frac{q-1}{2}}B^{q-1}-\beta^{\frac{q-1}{2}}A^{q-1}}$
$=\frac{1}{\alpha^{(\frac{q-1}{2})j}\beta^{(\frac{q-1}{2})j}}\sum_{A\in R_{1}}\frac{\beta^{\frac{q-1}{2}}}{\alpha^{\frac{q-1}{2}}}\frac{1}{A^{(q-1)j}}\frac{1}{\tilde{\pi}^{q-1}B^{q-1}-\beta^{q-1}A^{q-1}}$
$=-\frac{\beta^{(\frac{q-1}{2})j}}{\alpha^{(\frac{q-1}{2})j}}\frac{\beta^{\frac{q-1}{2}}}{\alpha^{\frac{q-1}{2}}}\sum_{A\in R_{1}}\frac{1}{(\beta A)^{(q-1)j}}\frac{1}{(\beta A)^{q-1}-({\tilde{\pi}B})^{q-1}}.$
@@
(2.10) $\quad\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}$
$=-\beta^{(\frac{q-1}{2})(j+1)}\sum_{A\in R_{1}}\frac{1}{(\beta A)^{(q-1)j}}\frac{1}{(\beta A)^{q-2}}\left(\frac{1}{\beta A}-\frac{1}{e(\beta A)}\right)$
$=-\beta^{(\frac{q-1}{2})(j+1)}\sum_{A\in R_{1}}\frac{1}{(\beta A)^{(q-1)j}}\frac{1}{(\beta A)^{q-2}}\left(\frac{1}{\beta A}-\frac{1}{e(\beta A)}\right)$
@@
$\qquad\qquad\qquad\qquad\qquad\quad-\left.\sum_{A\in R_{1}}\frac{1}{(\beta A)^{(q-1)j+(q-2)}e(\beta A)}\right).$
@@
(2.11) $\quad\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}$
@@
$\quad+\beta^{(\frac{q-1}{2})(j+1)}\sum_{A\in R_{1}}\frac{1}{(\beta A)^{(q-1)j+(q-2)}e(\beta A)}$
@@
$\quad+\beta^{-(\frac{q-1}{2})(j-1)-(q-2)}\sum_{A\in R_{1}}\frac{A^{-(q-1)j-(q-2)}}{e(\beta A)}.$
@@
(2.12) $\alpha^{-(\frac{q-1}{2})(j-1)-(q-2)}\sum_{B\in R_{1}}\frac{B^{-(q-1)j-(q-2)}}{e(\alpha B)}$
$\quad+\beta^{-(\frac{q-1}{2})(j-1)-(q-2)}\sum_{A\in R_{1}}\frac{A^{-(q-1)j-(q-2)}}{e(\beta A)}$

### v4

@@ 2.2. Hamahata’s Lambert series
$e_{C}(z)$ $=e_{\tilde{\pi}A}(z)$ (3.1)
$=z\prod_{0\neq\lambda\in\tilde{\pi}A}\left(1-\frac{z}{\lambda}\right)$
$=z\prod_{0\neq a\in A}\left(1-\frac{z}{\tilde{\pi}a}\right)$
$=z\prod_{a\in A_{+}}\left(1-\left(\frac{z}{\tilde{\pi}a}\right)^{q-1}\right),$
@@
$\frac{1}{e_{C}(z)}$ $=\frac{e_{C}^{\prime}(z)}{e_{C}(z)}$ (3.2)
$=\frac{1}{z}+\sum_{a\in A_{+}}\frac{-\frac{(q-1)z^{q-2}}{(\tilde{\pi}a)^{q-1}}}{1-\left(\frac{z}{\tilde{\pi}a}\right)^{q-1}}$
$=\frac{1}{z}-z^{q-2}\sum_{a\in A_{+}}\frac{1}{z^{q-1}-(\tilde{\pi}a)^{q-1}}.$
Thus for any $\alpha\in\Omega$ and $b\in A$, we have
$\frac{1}{e_{C}(\alpha b)}=\frac{1}{\alpha b}-(\alpha b)^{q-2}\sum_{a\in A_{+}}\frac{1}{(\alpha b)^{q-1}-(\tilde{\pi}a)^{q-1}}.$ (3.3)
@@
$\quad\alpha^{(\frac{q-1}{2})(j+1)}\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{(\alpha b)^{(q-1)j}\left((\alpha b)^{q-1}-(\tilde{\pi}a)^{q-1}\right)}$
$=\alpha^{(\frac{q-1}{2})(j+1)}\sum_{b\in A_{+}}\frac{1}{(\alpha b)^{(q-1)j}}\frac{1}{(\alpha b)^{q-2}}\left(\frac{1}{\alpha b}-\frac{1}{e_{C}(\alpha b)}\right)$
@@
$\left.\qquad\qquad\qquad\qquad\qquad\quad-\sum_{b\in A_{+}}\frac{1}{(\alpha b)^{(q-1)j+(q-2)}e_{C}(\alpha b)}\right).$
@@
$\quad\alpha^{(\frac{q-1}{2})(j+1)}\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{(\alpha b)^{(q-1)j}\left((\alpha b)^{q-1}-(\tilde{\pi}a)^{q-1}\right)}$ (3.4)
@@
$\quad-\alpha^{(\frac{q-1}{2})(j+1)}\sum_{b\in A_{+}}\frac{1}{(\alpha b)^{(q-1)j+(q-2)}e_{C}(\alpha b)}$
@@
$\quad\times\sum_{b\in A_{+}}\frac{b^{-(q-1)j-(q-2)}}{e_{C}(\alpha b)}.$
@@
$\quad\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{(\alpha b)^{(q-1)j}\left((\alpha b)^{q-1}-(\tilde{\pi}a)^{q-1}\right)}$
$=-\sum_{b\in A_{+}}\frac{1}{(\alpha b)^{(q-1)j}}\sum_{a\in A_{+}}\frac{1}{(\tilde{\pi}a)^{q-1}}$
$\quad+\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{(\alpha b)^{(q-1)(j-1)}(\tilde{\pi}a)^{q-1}\left((\alpha b)^{q-1}-(\tilde{\pi}a)^{q-1}\right)}$
$=-\sum_{b\in A_{+}}\frac{1}{(\alpha b)^{(q-1)j}}\sum_{a\in A_{+}}\frac{1}{(\tilde{\pi}a)^{q-1}}$
$\quad-\sum_{b\in A_{+}}\frac{1}{(\alpha b)^{(q-1)(j-1)}}\sum_{a\in A_{+}}\frac{1}{(\tilde{\pi}a)^{(q-1)2}}$
$\quad+\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{(\alpha b)^{(q-1)(j-2)}(\tilde{\pi}a)^{(q-1)2}\left((\alpha b)^{q-1}-(\tilde{\pi}a)^{q-1}\right)}$
$=\cdots$
$\quad(\textrm{inductively})$
$=-\sum_{b\in A_{+}}\frac{1}{(\alpha b)^{(q-1)j}}\sum_{a\in A_{+}}\frac{1}{(\tilde{\pi}a)^{q-1}}$
$\quad-\sum_{b\in A_{+}}\frac{1}{(\alpha b)^{(q-1)(j-1)}}\sum_{a\in A_{+}}\frac{1}{(\tilde{\pi}a)^{(q-1)2}}$
$\quad\cdots$
$\quad-\sum_{b\in A_{+}}\frac{1}{(\alpha b)^{q-1}}\sum_{a\in A_{+}}\frac{1}{(\tilde{\pi}a)^{(q-1)j}}$
$\quad+\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{(\tilde{\pi}a)^{(q-1)j}}\frac{1}{(\alpha b)^{q-1}-(\tilde{\pi}a)^{q-1}}$
$=-\sum_{k=1}^{j}\left(\sum_{b\in A_{+}}\frac{1}{(\alpha b)^{(q-1)(j-k+1)}}\sum_{a\in A_{+}}\frac{1}{(\tilde{\pi}a)^{(q-1)k}}\right)$
$\quad+\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{(\tilde{\pi}a)^{(q-1)j}}\frac{1}{(\alpha b)^{q-1}-(\tilde{\pi}a)^{q-1}}$
@@
$\quad+\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{(\tilde{\pi}a)^{(q-1)j}}\frac{1}{(\alpha b)^{q-1}-(\tilde{\pi}a)^{q-1}},$
@@
$\quad\alpha^{(\frac{q-1}{2})(j+1)}\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{(\alpha b)^{(q-1)(j+1)}\left((\alpha b)^{q-1}-(\tilde{\pi}a)^{q-1}\right)}$ (3.5)
@@
$\quad+\alpha^{(\frac{q-1}{2})(j+1)}\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{(\tilde{\pi}a)^{(q-1)j}}\frac{1}{(\alpha b)^{q-1}-(\tilde{\pi}a)^{q-1}}$
@@
$\quad+\alpha^{(\frac{q-1}{2})(j+1)}\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{(\tilde{\pi}a)^{(q-1)j}}\frac{1}{(\alpha b)^{q-1}-(\tilde{\pi}a)^{q-1}}.$
@@
$\quad-\alpha^{-(\frac{q-1}{2})(j-1)-(q-2)}\sum_{b\in A_{+}}\frac{b^{-(q-1)j-(q-2)}}{e_{C}(\alpha b)}$ (3.6)
@@
$\quad+\alpha^{(\frac{q-1}{2})(j+1)}\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{(\tilde{\pi}a)^{(q-1)j}}\frac{1}{(\alpha b)^{q-1}-(\tilde{\pi}a)^{q-1}}.$
@@
$\quad\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{(\tilde{\pi}a)^{(q-1)j}}\frac{1}{(\alpha b)^{q-1}-(\tilde{\pi}a)^{q-1}}$
$=\frac{1}{\tilde{\pi}^{(q-1)j}}\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{a^{(q-1)j}}\frac{1}{(\alpha b)^{q-1}-\alpha^{\frac{q-1}{2}}\beta^{\frac{q-1}{2}}a^{q-1}}$
$=\frac{1}{\tilde{\pi}^{(q-1)j}}\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{\alpha^{\frac{q-1}{2}}a^{(q-1)j}}\frac{1}{\alpha^{\frac{q-1}{2}}b^{q-1}-\beta^{\frac{q-1}{2}}a^{q-1}}$
$=\frac{1}{\alpha^{(\frac{q-1}{2})j}\beta^{(\frac{q-1}{2})j}}\sum_{a\in A_{+}}\frac{\beta^{\frac{q-1}{2}}}{\alpha^{\frac{q-1}{2}}}\frac{1}{a^{(q-1)j}}\frac{1}{\tilde{\pi}^{q-1}b^{q-1}-\beta^{q-1}a^{q-1}}$
$=-\frac{\beta^{(\frac{q-1}{2})j}}{\alpha^{(\frac{q-1}{2})j}}\frac{\beta^{\frac{q-1}{2}}}{\alpha^{\frac{q-1}{2}}}\sum_{a\in A_{+}}\frac{1}{(\beta a)^{(q-1)j}}\frac{1}{(\beta a)^{q-1}-({\tilde{\pi}b})^{q-1}}.$
@@
$\quad\alpha^{(\frac{q-1}{2})(j+1)}\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{(\tilde{\pi}a)^{(q-1)j}}\frac{1}{(\alpha b)^{q-1}-(\tilde{\pi}a)^{q-1}}$ (3.7)
$=-\beta^{(\frac{q-1}{2})(j+1)}\sum_{a\in A_{+}}\frac{1}{(\beta a)^{(q-1)j}}\frac{1}{(\beta a)^{q-2}}\left(\frac{1}{\beta a}-\frac{1}{e_{C}(\beta a)}\right)$
@@
$\qquad\qquad\qquad\qquad\qquad\quad-\left.\sum_{a\in A_{+}}\frac{1}{(\beta a)^{(q-1)j+(q-2)}e_{C}(\beta a)}\right).$
@@
$\quad\alpha^{(\frac{q-1}{2})(j+1)}\sum_{b\in A_{+}}\sum_{a\in A_{+}}\frac{1}{(\tilde{\pi}a)^{(q-1)j}}\frac{1}{(\alpha b)^{q-1}-(\tilde{\pi}a)^{q-1}}$ (3.8)
@@
$\quad+\beta^{(\frac{q-1}{2})(j+1)}\sum_{a\in A_{+}}\frac{1}{(\beta a)^{(q-1)j+(q-2)}e_{C}(\beta a)}$
@@
$\quad+\beta^{-(\frac{q-1}{2})(j-1)-(q-2)}\sum_{a\in A_{+}}\frac{a^{-(q-1)j-(q-2)}}{e_{C}(\beta a)}.$
@@
$\alpha^{-(\frac{q-1}{2})(j-1)-(q-2)}\sum_{b\in A_{+}}\frac{b^{-(q-1)j-(q-2)}}{e_{C}(\alpha b)}$
$\quad+\beta^{-(\frac{q-1}{2})(j-1)-(q-2)}\sum_{a\in A_{+}}\frac{a^{-(q-1)j-(q-2)}}{e_{C}(\beta a)}$
@@
Acknowledgements
The authors are enormously grateful to the anonymous referee for his/her careful reading of this paper, and for his/her insightful comments and suggestions.
Su Hu is supported by the Natural Science Foundation of Guangdong Province, China (No. 2024A1515012337). Min-Soo Kim is supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. NRF-2022R1F1A1065551 and No. RS-2025-23323090).
Statements and Declarations
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Authors declare that they do not have any conflict of interest.
Data availability
No data was used for the research described in the article.